Micheal buys two bags of chips and three boxes of pretzels for $5.13. He then buys another bag of chips and two more boxes of pretzels for $3.09. Find the cost of each bag of chips and each box of pretzels.
Answer by ad_alta(240)

Let 'p' be the price of pretzels and 'c' the price for chips. Then 2c+3p=5.13. Also, c+2p=3.09. The second equation tells us that c=3.09-2p. Substitute into the first equation and get 2(3.09-2p)+3p=5.13. This gives p=1.05 and c=0.99.